1) Use the Pythagorean Theorem to find the distance between X(7,11) and Y(-1,5)..Could you please show me your solution because I can't really understand this lesson!
2)Find the coordinates of the midpoint of a segment X(-4,3) and Y(-1,5)
3)Find the coordinates of A if B(0,5.5) is the midpoint of segment AC and C has coordinates (-3,6)
4) Find the coordinates of the missing endpoint given that S is the midpoint of segment RT: T(-4,3), S(-1,5)
Thank You very much=)

Answer by stanbon(75887)

1) Use the Pythagorean Theorem to find the distance between
X(7,11) and Y(-1,5)..
Could you please show me your solution because I can't really understand this lesson!
---
d^2 = (change in x)^2 + (change in y)^2
d^2 = [(7--1)^2 + (11-5)^2]
d^2 = [64+36]
d^2 = 100
distance = 10
----------------------------
2)Find the coordinates of the midpoint of a segment X(-4,3) and Y(-1,5)
x = (-4+-1)/2 = -5/2
y = (3+5)/2 = 4
midpoint::: (-5/2,4)
----------------------------
3)Find the coordinates of A if B(0,5.5) is the midpoint of segment AC and C has coordinates (-3,6)
Let the coordinate of A be (x,y)
(x+-3)/2 = 0
x-3 = 0
x = 3
-------
(y+6)/2 = 5.5
y+6 = 11
y = 5
Coordinates of A:: (3,5)
--------------------------------
4) Find the coordinates of the missing endpoint given that S is the midpoint of segment RT: T(-4,3), S(-1,5)
Let the coordinates of R are (x,y)
(x+-4)/2 = -1
x-4 = -2
x = 2
----
(y+3)/2 = 5
y+3 = 10
y = 7
-----
Endpoint: (2,7)
=====================
Cheers,
Stan H.
=====================